# Towards Understanding Why FixMatch Generalizes Better Than Supervised Learning

### v1

@@ Appendix A Theorem Statement
  • $N_{l}\geq N_{l,s}\cdot\mathrm{poly}(k)$, $\eta T\geq N_{l}\cdot\mathrm{poly}(k)$, and $\sqrt{d}\geq\eta T\cdot\mathrm{poly}(k)$.
@@
  • The number of unlabeled data for FixMatch $N_{c}\geq\eta T^{\prime}\cdot\mathrm{poly}(k)$ with $\eta T^{\prime}\geq\mathrm{poly}(k)$, and the ratio of single-view unlabeled data $\frac{N_{c,s}}{N_{c}}\leq\frac{k^{2}}{\eta sT^{\prime}}$.
@@
For sufficiently large $k>0$, for every $m=\mathrm{polylog}(k)$, $\eta\leq\frac{1}{\mathrm{poly}(k)}$, setting $T=\frac{\mathrm{poly}(k)}{\eta}$ and $T^{\prime}=\frac{\mathrm{poly}(k)}{\eta}$, when Parameter Assumption 10 is satisfied, with probability at least $1-e^{-\Omega(\log^{2}k)}$,
@@
    $\forall i\neq y:F_{y}^{(T+T^{\prime})}(X)\geq F_{i}^{(T+T^{\prime})}(X)+\Omega(\log k).$
  • (multi-view testing is good) for every $i,j\in[k]$, we have $\tilde{O}(1)\geq\Phi_{i}^{(T+T^{\prime})}\geq 0.4\Phi_{j}^{(T+T^{\prime})}+\Omega(\log k)$, and thus
    $\Pr_{(X,y)\in\mathcal{D}_{m}}\left[F_{y}^{(T+T^{\prime})}(X)\geq\max_{j\neq y}F_{j}^{(T+T^{\prime})}(X)+\Omega(\log k)\right]\geq 1-e^{-\Omega(\log^{2}k)}.$
  • (single-view testing is good) for every $i\in[k]$ and $l\in[2]$, we have $\Phi_{i,l}^{(T+T^{\prime})}\geq\Omega(\log k)$, and thus
    $\Pr_{(X,y)\in\mathcal{D}_{s}}\left[F_{y}^{(T+T^{\prime})}(X)\geq\max_{j\neq y}F_{j}^{(T+T^{\prime})}(X)+\Omega(\log k)\right]\geq 1-e^{-\Omega(\log^{2}k)}.$
For Semantic-Aware FixMatch (SA-FixMatch), we denote the number of unlabeled data in this case as $N_{u}$. Then we have the following assumption on $N_{u}$.
@@
Under Parameter Assumptions 10 and 12, SA-FixMatch achieves the same performance results as Theorem 11, but with a reduced requirement for the number of unlabeled data, decreasing from $N_{c}$ to $N_{u}$. Thus, we state the following theorem regarding the performance of SA-FixMatch.
@@
Under Parameter Assumption 10 and 12, we can achieve the same training and test performance as FixMatch in Theorem 11.
@@
Then in Phase II, the network predicts highly confident pseudo-labels for weakly-augmented samples and uses these correct pseudo-labels to supervise the corresponding strongly-augmented samples via consistency regularization. To theoretically analyze the learning process in Phase II, we build on the proof framework of Allen-Zhu & Li (2023) and demonstrate how the network learns the unlearned features while preserving the learned features during Phase II. Specifically, we present the induction hypothesis for Phase II in Appendix C, along with gradient calculations and function approximations for the unsupervised loss in Eq. (5) in Appendix D. We then provide a detailed proof of SA-FixMatch in Appendix E and extend the results to FixMatch in Appendix F. Finally, we generalize our proof to other FixMatch-like SSL methods in Appendix G.
Appendix B Results on Supervised Training
In this section, we first recall the results in supervised training that were derived in Allen-Zhu & Li (2023). Before showing their main results, we first introduce some necessary notations. For every $i\in[k]$, define $\Phi_{i,l}^{(t)}:=\sum_{r\in[m]}[\langle w_{i,r}^{(t)},v_{i,l}\rangle]^{+}$ and $\Phi_{i}^{(t)}:=\sum_{l\in[2]}\Phi_{i,l}^{(t)}$. Define
@@ Appendix B Results on Supervised Training
$\mathcal{M}_{F}:=\left\{(i,l^{*})\in[k]\times[2]\bigg{|}\Lambda_{i,l^{*}}^{(0)}\geq\Lambda_{i,3-l^{*}}^{(0)}\left(1+\frac{2}{\log^{2}m}\right)\right\}.$
The intuition behind $\mathcal{M}_{F}$ is that if $(i,l)\in\mathcal{M}_{F}$, then the feature $v_{i,l}$ is likely to be learned by the model and the feature $v_{i,3-l}$ is likely to be missed due to model initialization and data distribution. $\mathcal{M}_{F}$ satisfies the following property (refer to the Proposition C.2. of Allen-Zhu & Li (2023)):
@@
Suppose $m\leq\mathrm{poly}(k)$. For every $i\in[k]$, $\Pr[(i,1)\in\mathcal{M}_{F}~{}\mbox{or }(i,2)\in\mathcal{M}_{F}]\geq 1-o(1)$.
Based on Theorem 1 of Allen-Zhu & Li (2023), after training for $T$ iterations with the supervised training loss $L_{s}^{(t)}$
$L_{s}^{(t)}=\mathbb{E}_{(X,y)\sim\mathcal{Z}_{l}}\left[-\log\mathbf{logit}_{y}(F^{(t)},X)\right],$
the training accuracy on labeled samples is perfect and $L_{s}^{(T)}$ approaches zero, i.e., for every $(X,y)\in\mathcal{Z}_{l}$,
@@
This means that the test accuracy of multi-view data is good. However, for the single-view data $(X,y)\sim\mathcal{D}_{s}$, whenever $(i,l^{*})\in\mathcal{M}_{F}$, we have $\Phi_{i,3-l^{*}}^{(T)}\ll\frac{1}{\mathrm{polylog}(k)}$ and
@@
which means that the test accuracy on single-view data is nearly $50\%$.
The results in Allen-Zhu & Li (2023) fully indicate the feature learning process of supervised learning. The main reason for the imperfect performance of SL is that, due to "lottery winning", it only captures one of the two features for each semantic class during the supervised training process. Therefore for single-view data without this feature, it has low test accuracy.
@@
where $b=\operatorname*{arg\,max}_{i\in[k]}\mathbf{logit}_{i}(F^{(t)},\alpha(X))$, $\tau$ is the confidence threshold and $\alpha,\mathcal{A}$ are the weak and strong augmentations, respectively. For the simplicity of proof, we set $\alpha$ to be identity mapping. In the following, we will prove Theorem 11. By setting $\tau=1-\tilde{O}(1/s^{2})$, we will show that after we train the network $F^{(T)}$ combining the loss $L_{u}^{(t)}$ for another $T^{\prime}=\frac{\mathrm{poly}(k)}{\eta}$ epochs, the network can learn complete semantic features for all classes, resulting in perfect test performance on both multi-view and single-view data.
@@ Appendix C Induction Hypothesis
In this section, to prove our theorem, similar to Allen-Zhu & Li (2023), we present an induction hypothesis for every training iteration $t$ in learning Phase II. We first show the loss function in learning Phase II.
@@
In Phase II, after we train the network $F$ for $T=\frac{\mathrm{poly}(k)}{\eta}$ epochs using $L_{s}^{(t)}$ in the Phase I, according to the results in Appendix B, one of the features in each class is captured. Then we consider to optimize the network $F^{(T)}$ using the following combination of losses:
@@
where $b=\operatorname*{arg\,max}_{i\in[k]}\mathbf{logit}_{i}(F^{(t)},\alpha(X))$. Recall $\tau=1-\tilde{O}(1/s^{2})$, when $t\geq T$ and we use $F^{(t)}$ to classify the unlabeled data $X\sim\mathcal{Z}_{u}$, we will get a correct pseudo-label with high probability, i.e., $b=y$, where $y$ denotes the ground truth label of $X$. This means that for $X\sim\mathcal{Z}_{u,m}$, with probability at least $1-e^{-\Omega(\log^{2}k)}$, $\mathbf{logit}_{y}(F^{(t)},X)\geq\tau$ and for $X\sim\mathcal{Z}_{u,s}$, when $(y,l^{*})\in\mathcal{M}_{F}$ and $\hat{l}(X)=l^{*}$, with the probability at least $1-e^{-\Omega(\log^{2}k)}$, $\mathbf{logit}_{y}(F^{(t)},X)\geq\tau$. We denote the samples in $\mathcal{Z}_{u}$ that satisfy $\mathbf{logit}_{y}(F^{(t)},X)\geq\tau$ as $\tilde{\mathcal{Z}}_{u}$ and let $\tilde{N}_{u}=|\tilde{\mathcal{Z}}_{u}|$. In this way, we can further simplify the loss as
@@
For every $l\in[2]$, for every $r\in[m]$, for every $X\in\tilde{\mathcal{Z}}_{u}$ and $i\in[k]$,
@@
    $\Phi_{i,l}^{(t)}\geq\Omega(\sigma_{0})$ and $\Phi_{i,l}^{(t)}\leq\tilde{O}(1)$.
@@
The intuition behind Induction Hypothesis 15 is that training with semi-supervised loss (11) filters out feature noises and background noises for both multi-view data and single-view data. This can be seen in comparison with Induction Hypothesis C.3 of Allen-Zhu & Li (2023). With the help of Induction Hypothesis 15, we can prove that the correlations between $w_{i,r}$ and learned features in Phase I are retained in Phase II, and the correlations between $w_{i,r}$ and unlearned features will increase to a large scale ($\log(k)$) in the end of learning Phase II.
@@ Appendix E Proof for Semantic-Aware FixMatch
For theoretical proof, we assume that Grad-CAM in SA-CutOut can correctly identify the learned feature after the first stage. In this case, the formulation of strong augmentation $\mathcal{A}(\cdot)$ with SA-CutOut for $X\sim\tilde{\mathcal{Z}}_{u}$ and $(i,l^{*})\in\mathcal{M}_{F}\cap\mathcal{V}(X)$ ($l^{*}$ varies depending on $i$) is
@@ E.1 Useful Claims
Based on the results from Allen-Zhu & Li (2023), at the end of learning Phase I, for $(i,l^{*})\in\mathcal{M}_{F}$, we have $\Phi_{i,l^{*}}^{(T)}\geq\Omega(\log k)$ while $\Phi_{i,3-l^{*}}^{(T)}\leq 1/\mathrm{polylog}(k)$. Below the first claim addresses the initial growth of the correlations between $w_{i,r}$ and the unlearned feature ($\Phi_{i,3-l^{*}}^{(t)}$) in learning Phase II, and the second claim asserts that the correlations between $w_{i,r}$ and the learned feature ($\Phi_{i,l^{*}}^{(t)}$) are preserved during learning Phase II. Here we first give a naive bound on the $\mathbf{logit}$ function based on function approximation result Claim 21.
@@
  • for every $(X,y)\sim\tilde{\mathcal{Z}}_{u,m}$ and $(i,l^{*})\in\mathcal{M}_{F}$: $\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X))=O\left(\frac{e^{O(\Phi_{i,3-l^{*}})}}{e^{O(\Phi_{i,3-l^{*}})}+k}\right)$.
@@
Recall $F_{i}^{(t)}(\mathcal{A}(X))=\sum_{r\in[m]}\sum_{p\in[P]}\overline{\operatorname{ReLU}}(\langle w_{i,r},\mathcal{A}(x_{p})\rangle)$. According to Claim 21, data assumption 7 and data augmentation defined in (14), we have that for $(X,y)\sim\tilde{\mathcal{Z}}_{u,m}$ and $(i,l^{*})\in\mathcal{M}_{F}$ ($l^{*}$ varies depending on $i$),
@@
Thus, combining the above results, for every $(X,y)\sim\tilde{\mathcal{Z}}_{u,m}$ and $(i,l^{*})\in\mathcal{M}_{F}$, we have for every $i\in[k]$,
@@
On the other hand, for the single-view data in $\tilde{\mathcal{Z}}_{u,s}$, as the only class-specific semantic feature is masked after we conduct strong augmentation, only noisy unlearned features and background noises remain. Thus, for $(X,y)\sim\tilde{\mathcal{Z}}_{u,s}$ and $(i,l^{*})\in\mathcal{M}_{F}$, we have for $i\neq y$,
@@
Now we are ready to prove the following claim on the initial growth of $\Phi_{i,3-l^{*}}^{(t)}$ with $(i,l^{*})\in\mathcal{M}_{F}$.
@@
Suppose Induction Hypothesis 15 holds at iteration $t$, then for every $i\in[k]$ with $(i,l^{*})\in\mathcal{M}_{F}$, suppose $\Phi_{i,3-l^{*}}^{(t)}\leq O(1)$, then it satisfies
@@
Based on the results from Allen-Zhu & Li (2023), when $t\geq T$, we have
@@
i.e., the supervised loss is fully minimized in the first stage and contributes little in the second stage.
@@
Consider $r=\operatorname*{arg\,max}_{r\in[m]}\{\langle w_{i,r}^{(t)},v_{i,3-l^{*}}\rangle\}$, then as $m=\mathrm{polylog}(k)$, we know $\langle w_{i,r}^{(t)},v_{i,3-l^{*}}\rangle\geq\tilde{\Omega}(\Phi_{i,3-l^{*}}^{(t)})$. Recall $V_{i,r,3-l^{*}}(X):=\mathbb{I}_{v_{i,3-l^{*}}\in\mathcal{V}(X)}\sum_{p\in\mathcal{P}_{v_{i,3-l^{*}}}(X)}\overline{\operatorname{ReLU}}^{\prime}(\langle w_{i,r}^{(t)},\mathcal{A}(x_{p})\rangle)\mathcal{A}(z_{p})$, according to Induction Hypothesis 15(a) and definition in (14), we have
@@
Define iteration threshold $T_{0}:=\tilde{\Theta}\left(\frac{k}{\eta\sigma_{0}^{q-2}}\right)$, then for every $i\in[k],(i,l^{*})\in\mathcal{M}_{F}$ and $t\geq T+T_{0}$, it satisfies that $\Phi_{i,3-l^{*}}^{(t)}=\Theta(1)$.
As we stated in Claim 23, model prediction for the augmented single-view data in $\tilde{\mathcal{Z}}_{u,s}$ are kept to the scale of $O\left(\frac{1}{k}\right)$, since after strong augmentation there remains only noises. Now we present the convergence of multi-view data in $\tilde{\mathcal{Z}}_{u,m}$ from $T+T_{0}$ till the end.
@@
Suppose that the Induction Hypothesis 15 holds for every iteration $T<t\leq T+T^{\prime}$, and suppose $\frac{\tilde{N}_{u,s}}{\tilde{N}_{u}}\leq\frac{k^{2}}{\eta sT^{\prime}}$, then
$\sum_{t=T+T_{0}}^{T+T^{\prime}}\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}[1-\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X))]\leq\tilde{O}\left(\frac{k}{\eta}\right).$
@@
Take $r=\operatorname*{arg\,max}_{r\in[m]}\{\langle w_{i,r}^{(t)},v_{i,3-l^{*}}\rangle\}$, then by $m=\mathrm{polylog}(k)$ we know $\langle w_{i,r}^{(t)},v_{i,3-l^{*}}\rangle\geq\tilde{\Omega}(\Phi_{i,3-l^{*}}^{(t)})=\tilde{\Omega}(1)$ for $t\geq T+T_{0}$.

### v2

@@ 6 Conclusion
Acknowledgments and Author Contributions
J. Pan and V. Y. F. Tan are supported by a Singapore Ministry of Education (MOE) AcRF Tier 2 grant (A-8000423-00-00). J. Li led the development of the theoretical framework, provided the rigorous proofs, and also led the experimental design and implementation. J. Pan contributed the initial sketch of the main proof. All authors discussed the results and contributed to the final manuscript.
@@ Appendix A Theorem Statement
  • $N_{l}\geq N_{l,s}\cdot\mathrm{poly}(k)$, $\eta T_{1}\geq N_{l}\cdot\mathrm{poly}(k)$, and $\sqrt{d}\geq\eta T_{1}\cdot\mathrm{poly}(k)$.
@@
  • The number of unlabeled data for FixMatch $N_{u}\geq\eta T_{2}\cdot\mathrm{poly}(k)$ with $\eta T_{2}\geq\mathrm{poly}(k)$, and the ratio of single-view unlabeled data $\frac{N_{u,s}}{N_{u}}\leq\frac{k^{2}}{\eta sT_{2}}$.
@@
For sufficiently large $k>0$, for every $m=\mathrm{polylog}(k)$, $\eta\leq\frac{1}{\mathrm{poly}(k)}$, setting $T=T_{1}+T_{2}$ with $T_{1}=\frac{\mathrm{poly}(k)}{\eta}$ and $T_{2}=\frac{\mathrm{poly}(k)}{\eta}$, when Parameter Assumption 10 is satisfied, with probability at least $1-e^{-\Omega(\log^{2}k)}$,
@@
    $\forall i\neq y:F_{y}^{(T)}(X)\geq F_{i}^{(T)}(X)+\Omega(\log k).$
  • (multi-view testing is good) for every $i,j\in[k]$, we have $\tilde{O}(1)\geq\Phi_{i}^{(T)}\geq 0.4\Phi_{j}^{(T)}+\Omega(\log k)$, and thus
    $\Pr_{(X,y)\in\mathcal{D}_{m}}\left[F_{y}^{(T)}(X)\geq\max_{j\neq y}F_{j}^{(T)}(X)+\Omega(\log k)\right]\geq 1-e^{-\Omega(\log^{2}k)}.$
  • (single-view testing is good) for every $i\in[k]$ and $l\in[2]$, we have $\Phi_{i,l}^{(T)}\geq\Omega(\log k)$, and thus
    $\Pr_{(X,y)\in\mathcal{D}_{s}}\left[F_{y}^{(T)}(X)\geq\max_{j\neq y}F_{j}^{(T)}(X)+\Omega(\log k)\right]\geq 1-e^{-\Omega(\log^{2}k)}.$
For Semantic-Aware FixMatch (SA-FixMatch), we denote the number of unlabeled data in this case as $N_{c}$. Then we have the following assumption on $N_{c}$.
@@
Under Parameter Assumptions 10 and 12, SA-FixMatch achieves the same performance results as Theorem 11, but with a reduced requirement for the number of unlabeled data, decreasing from $N_{u}$ to $N_{c}$. Thus, we state the following theorem regarding the performance of SA-FixMatch.
@@
Under Parameter Assumption 10 and 12, SA-FixMatch can achieve the same training and test performance as FixMatch in Theorem 11.
@@
Then in Phase II, the network predicts highly confident pseudo-labels for weakly-augmented samples and uses these correct pseudo-labels to supervise the corresponding strongly-augmented samples via consistency regularization. To theoretically analyze the learning process in Phase II, we build on the proof framework of Allen-Zhu & Li (2023) and demonstrate how the network learns the unlearned features while preserving the learned features during Phase II. Specifically, we present the induction hypothesis for Phase II in Appendix C, along with gradient calculations and function approximations for the unsupervised loss Eq. (5) in Appendix D. We then provide a detailed proof of SA-FixMatch in Appendix E and extend the results to FixMatch in Appendix F. Finally, we generalize our proof to other FixMatch-like SSL methods in Appendix G.
Appendix B Results on Supervised Learning
In this section, we first recall the results in SL that were derived in Allen-Zhu & Li (2023). Before showing their main results, we first introduce some necessary notations. For every $i\in[k]$, define $\Phi_{i,l}^{(t)}:=\sum_{r\in[m]}[\langle w_{i,r}^{(t)},v_{i,l}\rangle]^{+}$ and $\Phi_{i}^{(t)}:=\sum_{l\in[2]}\Phi_{i,l}^{(t)}$. Define
@@ Appendix B Results on Supervised Learning
:=$\mathcal{M}:=\left\{(i,l^{*})\in[k]\times[2]\bigg{|}\Lambda_{i,l^{*}}^{(0)}\geq\Lambda_{i,3-l^{*}}^{(0)}\left(1+\frac{2}{\log^{2}m}\right)\right\}.$
The intuition behind $\mathcal{M}$ is that, subject to model initialization, if $(i,l)\in\mathcal{M}$, then the feature $v_{i,l}$ will be learned by the model during supervised learning process and the feature $v_{i,3-l}$ will be missed. The set $\mathcal{M}$ satisfies the following property (refer to the Proposition C.2. of Allen-Zhu & Li (2023)):
@@
Suppose $m\leq\mathrm{poly}(k)$. For every $i\in[k]$, $\Pr[(i,1)\in\mathcal{M}~{}\mbox{or }(i,2)\in\mathcal{M}]\geq 1-o(1)$.
Based on Theorem 1 of Allen-Zhu & Li (2023), after training for $T$ iterations with the supervised training loss $L_{s}^{(t)}=\mathbb{E}_{(X,y)\sim\mathcal{Z}_{l}}\left[-\log\mathbf{logit}_{y}(F^{(t)},X)\right],$ the training accuracy on labeled samples is perfect and $L_{s}^{(T)}$ approaches zero, i.e., for every $(X,y)\in\mathcal{Z}_{l}$,
@@
This means that the test accuracy of multi-view data is good. However, for the single-view data $(X,y)\sim\mathcal{D}_{s}$, whenever $(i,l^{*})\in\mathcal{M}$, we have $\Phi_{i,3-l^{*}}^{(T)}\ll\frac{1}{\mathrm{polylog}(k)}$ and
@@
which means that the test accuracy on single-view data is nearly 50%.
The results in Allen-Zhu & Li (2023) fully indicate the feature learning process of SL. The main reason for the imperfect performance of SL is that, due to "lottery winning", it only captures one of the two semantic features for each class during the supervised training process. Therefore, for single-view data without this feature, it has low test accuracy.
@@
where $b=\operatorname*{arg\,max}_{i\in[k]}\mathbf{logit}_{i}(F^{(t)},\alpha(X))$, $\tau$ is the confidence threshold and $\alpha,\mathcal{A}$ are the weak and strong augmentations, respectively. For the simplicity of proof, we set $\alpha$ to be identity mapping. In the following, we will prove Theorem 11. By setting $\tau=1-\tilde{O}(1/s^{2})$, we will show that after training the supervised network $F^{(T_{1})}$ with the unsupervised loss $L_{u}^{(t)}$ for an additional $T_{2}=\frac{\mathrm{poly}(k)}{\eta}$ epochs, the FixMatch-trained network $F^{(T)}$ learns complete semantic features for all classes, achieving perfect test performance on both multi-view and single-view data.
@@ Appendix C Induction Hypothesis
In this section, to prove our theorem, similar to Allen-Zhu & Li (2023), we present an induction hypothesis for every training iteration $t$ in Phase II. We first show the loss function in Phase II.
@@
In Phase II, after we train the network $F$ for $T_{1}=\frac{\mathrm{poly}(k)}{\eta}$ epochs using $L_{s}^{(t)}$ in the Phase I, according to the results in Appendix B, one of the features in each class is captured. Then we consider to optimize the network $F^{(T_{1})}$ using the following combination of losses:
@@
where $b=\operatorname*{arg\,max}_{i\in[k]}\mathbf{logit}_{i}(F^{(t)},\alpha(X))$. Recall $\tau=1-\tilde{O}(1/s^{2})$, when $t\geq T_{1}$ and we use $F^{(t)}$ to classify the unlabeled data $X\sim\mathcal{Z}_{u}$, we will get a correct pseudo-label with high probability, i.e., $b=y$, where $y$ denotes the ground truth label of $X$. This means that for $X\sim\mathcal{Z}_{u,m}$, with probability at least $1-e^{-\Omega(\log^{2}k)}$, $\mathbf{logit}_{y}(F^{(t)},X)\geq\tau$ and for $X\sim\mathcal{Z}_{u,s}$, when $(y,l^{*})\in\mathcal{M}$ and $\hat{l}(X)=l^{*}$, with the probability at least $1-e^{-\Omega(\log^{2}k)}$, $\mathbf{logit}_{y}(F^{(t)},X)\geq\tau$. We denote the samples in $\mathcal{Z}_{u}$ that satisfy $\mathbf{logit}_{y}(F^{(t)},X)\geq\tau$ as $\tilde{\mathcal{Z}}_{u}$ and let $\tilde{N}_{u}=|\tilde{\mathcal{Z}}_{u}|$. In this way, we can further simplify the loss as
@@
During Phase II ($t\geq T_{1}$), for every $l\in[2]$, for every $r\in[m]$, for every $X\in\tilde{\mathcal{Z}}_{u}$ and $i\in[k]$,
@@
    $\Phi_{i,l}^{(t)}\geq\tilde{\Omega}(\sigma_{0})$ and $\Phi_{i,l}^{(t)}\leq\tilde{O}(1)$.
@@
The intuition behind Induction Hypothesis 15 is that training with semi-supervised loss Eq. (11) filters out feature noises and background noises for both multi-view data and single-view data. This can be seen in comparison with Induction Hypothesis C.3 of Allen-Zhu & Li (2023). With the help of Induction Hypothesis 15, we can prove that the correlations between $w_{i,r}$ and learned features in Phase I are retained in Phase II, and the correlations between $w_{i,r}$ and unlearned features will increase to a large scale ($\log(k)$) in the end of learning Phase II.
@@ Appendix E Proof for Semantic-Aware FixMatch
For theoretical proof, we assume that Grad-CAM in SA-CutOut can correctly identify the learned feature after the first stage. In this case, the formulation of strong augmentation $\mathcal{A}(\cdot)$ with SA-CutOut for $X\sim\tilde{\mathcal{Z}}_{u}$ and $(i,l^{*})\in\mathcal{M}\cap\mathcal{V}(X)$ ($l^{*}$ varies depending on $i$) is
@@ E.1 Useful Claims
Based on the results from Allen-Zhu & Li (2023), at the end of learning Phase I, for $(i,l^{*})\in\mathcal{M}$, we have $\Phi_{i,l^{*}}^{(T_{1})}\geq\Omega(\log k)$ while $\Phi_{i,3-l^{*}}^{(T_{1})}\leq 1/\mathrm{polylog}(k)$. Below the first claim addresses the initial growth of the correlations between $w_{i,r}$ and the unlearned feature ($\Phi_{i,3-l^{*}}^{(t)}$) in learning Phase II, and the second claim asserts that the correlations between $w_{i,r}$ and the learned feature ($\Phi_{i,l^{*}}^{(t)}$) are preserved during learning Phase II. Here we first give a naive bound on the $\mathbf{logit}$ function based on function approximation result Claim 21.
@@
  • for every $(X,y)\sim\tilde{\mathcal{Z}}_{u,m}$ and $(i,l^{*})\in\mathcal{M}$: $\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X))=O\left(\frac{e^{O(\Phi_{i,3-l^{*}})}}{e^{O(\Phi_{i,3-l^{*}})}+k}\right)$.
@@
Recall $F_{i}^{(t)}(\mathcal{A}(X))=\sum_{r\in[m]}\sum_{p\in[P]}\overline{\operatorname{ReLU}}(\langle w_{i,r},\mathcal{A}(x_{p})\rangle)$. According to Claim 21, data distribution Def. 7 and data augmentation defined in Eq. (14), we have that for $(X,y)\sim\tilde{\mathcal{Z}}_{u,m}$ and $(i,l^{*})\in\mathcal{M}$ ($l^{*}$ varies depending on $i$),
@@
Thus, combining the above results, for every $(X,y)\sim\tilde{\mathcal{Z}}_{u,m}$ and $(i,l^{*})\in\mathcal{M}$, we have for every $i\in[k]$,
@@
On the other hand, for the single-view data in $\tilde{\mathcal{Z}}_{u,s}$, as the only class-specific semantic feature is removed after we conduct strong augmentation, only noisy unlearned features and background noises remain. Thus, for $(X,y)\sim\tilde{\mathcal{Z}}_{u,s}$ and $(i,l^{*})\in\mathcal{M}$, we have for $i\neq y$,
@@
Now we are ready to prove the following claim on the initial growth of $\Phi_{i,3-l^{*}}^{(t)}$ with $(i,l^{*})\in\mathcal{M}$.
@@
Suppose Induction Hypothesis 15 holds at iteration $t$, then for every $i\in[k]$ with $(i,l^{*})\in\mathcal{M}$, suppose $\Phi_{i,3-l^{*}}^{(t)}\leq O(1)$, then it satisfies
@@
Based on the results from Allen-Zhu & Li (2023), when $t\geq T_{1}$, we have
@@
i.e., the supervised loss is fully minimized in Phase I and contributes little in Phase II.
@@
Consider $r=\operatorname*{arg\,max}_{r\in[m]}\{\langle w_{i,r}^{(t)},v_{i,3-l^{*}}\rangle\}$, then as $m=\mathrm{polylog}(k)$, we know $\langle w_{i,r}^{(t)},v_{i,3-l^{*}}\rangle\geq\tilde{\Omega}(\Phi_{i,3-l^{*}}^{(t)})$. Recall $V_{i,r,3-l^{*}}(X):=\mathbb{I}_{v_{i,3-l^{*}}\in\mathcal{V}(X)}\sum_{p\in\mathcal{P}_{v_{i,3-l^{*}}}(X)}\overline{\operatorname{ReLU}}^{\prime}(\langle w_{i,r}^{(t)},\mathcal{A}(x_{p})\rangle)\mathcal{A}(z_{p})$, according to Induction Hypothesis 15(a) and definition in Eq. (14), we have
@@
Define iteration threshold $T_{0}:=\tilde{\Theta}\left(\frac{k}{\eta\sigma_{0}^{q-2}}\right)$, then for every $i\in[k],(i,l^{*})\in\mathcal{M}$ and $t\geq T_{1}+T_{0}$, it satisfies that $\Phi_{i,3-l^{*}}^{(t)}=\Theta(1)$.
As we stated in Claim 23, model prediction for the augmented single-view data in $\tilde{\mathcal{Z}}_{u,s}$ are kept to the scale of $O\left(\frac{1}{k}\right)$, since after strong augmentation there remains only noises. Now we present the convergence of multi-view data in $\tilde{\mathcal{Z}}_{u,m}$ from $T_{1}+T_{0}$ till the end.
@@
Suppose that the Induction Hypothesis 15 holds for every iteration $T_{1}<t\leq T_{1}+T_{2}$, and suppose $\frac{\tilde{N}_{c,s}}{\tilde{N}_{c}}\leq\frac{k^{2}}{\eta sT_{2}}$, then
$\sum_{t=T_{1}+T_{0}}^{T_{1}+T_{2}}\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}[1-\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X))]\leq\tilde{O}\left(\frac{k}{\eta}\right).$
@@
Take $r=\operatorname*{arg\,max}_{r\in[m]}\{\langle w_{i,r}^{(t)},v_{i,3-l^{*}}\rangle\}$, then by $m=\mathrm{polylog}(k)$ we know $\langle w_{i,r}^{(t)},v_{i,3-l^{*}}\rangle\geq\tilde{\Omega}(\Phi_{i,3-l^{*}}^{(t)})=\tilde{\Omega}(1)$ for $t\geq T_{1}+T_{0}$.